OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. −6x + 3y = 18 4x − 3y = 6 2x + 4y = 24 6x + 12y = 36 3x − y = 14 −9x + 3y = −42 5x + 2y = 13 −x + 4y = −6

OpenStudy (anonymous):

Please help me

OpenStudy (welshfella):

infinitely many solutions occur when one equation is a multiple of the other. Basically they produce they produce the same line in the coordinate plane.

OpenStudy (anonymous):

Okay.. so what would I do to find the answer?

OpenStudy (welshfella):

an example would be x + y = 6 -2x - 2y = -12 - because the second equation = first multiplied by -2

OpenStudy (welshfella):

see if you can find something similar in the 4 given options are the y , x terms and the number in one equation multiplied by the same value to get the other?

OpenStudy (anonymous):

Oh okay, I'll look back at them and see. Thank you so much !!

OpenStudy (welshfella):

yw Hint - at first it looks like number 2 is the one because 2x* 3 = 6x and 4y * 3 = 12y but 24 *3 does not equal 36 so its not that one.

OpenStudy (anonymous):

Yeahh true ... hm. Algebra is my hardest so I dont really know what to do exactly

OpenStudy (welshfella):

look at number 3 how do you go from 3x to -9x, and -y to 3y and 14 to -42?

OpenStudy (anonymous):

It multiplies by 3 ?

OpenStudy (anonymous):

-3*

OpenStudy (welshfella):